A car radiator has a capacity of 16qts and is filled w/a 25% antifreeze solution. How much must be drained off and replaced w/pure antifreeze to obtain a 40% antifreeze solution?
Answer by ankor@dixie-net.com(22740)

A car radiator has a capacity of 16 qts and is filled w/a 25% anti-freeze solution. How much must be drained off and replaced w/pure antifreeze to obtain a 40% anti-freeze solution?
:
Let x = the amt drained and the amt added
:
Drained + Added = required mixture
.25(16-x) + 1.0x = .40(16)
:
4 - .25x + 1x = .40(16)
:
4 + .75x = 6.4
:
.75x = 6.4 - 4
:
.75x = 2.4
:
x = 2.4/.75
:
x = 3.2 gal of pure antifreeze
:
:
Check solution using x = 3.2:
.25(16-3.2) + 3.2 = .40(16)
.25(12.8) + 3.2 = 6.4
3.2 + 3.2 = 6.4